OpenStudy (anonymous):

First order differential equations-- Suppose there's a lake. A population of fish P(t) is attacked by a disease in time t=0; resulting in fishes stopping their reproduction (birth rate = 0), and a death rate proportional to 1/(sqrt(P)). If there were initially 900 fish in the tank and 441 were lost after 6 weeks, how long will it take for all fishes to die?

OpenStudy (anonymous):

According to the book, the answer is 20 weeks, or t=20, i'm thinking I need to integrate there?

OpenStudy (phi):

you want to put the p term with the dp term and the dt on the other side. then integrate

OpenStudy (anonymous):

Having only one constant, how can I evaluate for two different points of time? 1st should be t(0) = 900, and the second t(6) = 459?

OpenStudy (anonymous):

Solving for C gives me 900 for the initial population in time = 0, but not for the 459 in t (6). Also not for t = 20 weeks.

OpenStudy (phi):

I will post something if I can make sense of it...

OpenStudy (phi):

The death rate is proportional to 1/sqrt(P). The number of fish dying is P times the death rate or \[ \frac{dP}{dt} = -C_0P\cdot P^{-\frac{1}{2}}\] \[ \frac{dP}{dt} = P^{\frac{1}{2}}\] separate the variables: \[ P^{-\frac{1}{2}} dP= -C_0 dt\] integrate to get \[ 2\sqrt{P}= -C_0 t + C_1 \] as these are unknown constants, we can divide by 2, and use different constants \[\sqrt{P}= -At+B \] at t=0, P=900 and we find B=30 now, if we use 441 as the number still alive (because 441 is a perfect square), and solve for A at t=6: \[ \sqrt{441}= -6A+30 \] \[ A= (21-30)/-6= \frac{3}{2} \] \[ \sqrt{P}= -1.5t+30 \] setting P = 0 (no more fish!) \[ t= 30\cdot \frac{2}{3}= 20 \]

OpenStudy (anonymous):

Thank you SO MUCH!, I was seriously stuck at the two constants part. I knew there had to be two, I just didn't know where I could stick them in into the problem. Thanks lots!